OpenStudy (anonymous):

Randy drives his car to work each day. Suppose the expression 6/x + 6/3x represents the total time it takes Randy to drive to work in a day, where x represents the average speed of the car in miles per hour, during the first part of the drive. Use addition to simplify the expression

OpenStudy (anonymous):

My answer is 4/9 - is that correct?

OpenStudy (bohotness):

how do you come up with \(\4\9\)

OpenStudy (anonymous):

6/x * 3/3 = 18/3x + 6/3x 18 + 6 =24 24/3x

OpenStudy (anonymous):

3 * 18 = 54 24/54 = 4/9

OpenStudy (bohotness):

okay

OpenStudy (anonymous):

Did I do it correctly?

OpenStudy (bohotness):

i think you did but i am not sure

OpenStudy (anonymous):

Any help from anyone else? Please?

OpenStudy (perl):

be right back *